Where must a 500 N weight be hang on a uniform beam with a length of 5 m and weighing 150 N so that a woman at one end of the beam carries only one fourth as much load as a man at the other end?
1
Expert's answer
2021-08-09T08:03:49-0400

"F_{woman}= \frac{1}{4}* F_{man}\\
\sum F_y = 0\\
500+150= F_{woman}+F_{man}\\
650=F_{woman}+4F_{woman}\\
4F_{woman}=650\\
F_{woman}130 N\\
\sum M_{man}= 0\\
F_{woman} * l -150 * \frac{l}{2}-500 * x =0 \\
130*5 -150 * \frac{5}{2}-500*x =0\\
x= 0.55 m"

So, the weight needed should be suited at 0.55 m away from the man.